OpenStudy (anonymous):

What is the midpoint of the line segment with endpoints (5, -2) and (1, -6)? Select one of the options below as your answer: A. (2, -2) . (3, -4) C. (3, -6) D. (4, -3)

OpenStudy (lgbasallote):

\[\large midpoint = (\frac{x_1 + x_2}{2} , \frac{y_1 + y_2}{2})\]

OpenStudy (lgbasallote):

hehe :p

sam (.sam.):

\[\Huge M=(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2})\] \[\Huge M=(\frac{5+1}{2},\frac{-2+-6}{2})\]

OpenStudy (anonymous):

3,-4?

OpenStudy (lgbasallote):

sounds right to me

OpenStudy (lgbasallote):

dont forget the parentheses though